# Learning Robust Social Strategies with Large Language Models

### v1

@@ Abstract
As agentic AI becomes more widespread, agents with distinct and possibly conflicting goals will interact in complex ways. These multi-agent interactions pose a fundamental challenge, particularly in social dilemmas, where agents’ individual incentives can undermine collective welfare. While reinforcement learning (RL) has been effective for aligning large language models (LLMs) in the single-agent regime, prior small-network results suggest that standard RL in multi-agent settings often converges to defecting, self-interested policies. We show the same effect in LLMs: despite cooperative priors, RL-trained LLM agents develop opportunistic behavior that can exploit even advanced closed-source models. To address this tendency of RL to converge to poor equilibria, we adapt a recent opponent-learning awareness algorithm, Advantage Alignment, to fine-tune LLMs toward multi-agent cooperation and non-exploitability. We then introduce a group-relative baseline that simplifies advantage computation in iterated games, enabling multi-agent training at LLM scale. We also contribute a novel social dilemma environment, Trust and Split, which requires natural language communication to achieve high collective welfare. Across a wide range of social dilemmas, policies learned with Advantage Alignment achieve higher collective payoffs while remaining robust against exploitation by greedy agents.
@@ 1 Introduction
LLMs undergo large-scale pretraining, instruction tuning, and reinforcement learning, and continue to exhibit increasingly advanced capabilities (deepseekr1). Coupled with decreasing deployment costs and improved adaptability to downstream tasks, these trends enhance the commercial and practical viability of LLM agents across a wide range of applications. Recent efforts are already translating this potential into concrete systems. Anthropic’s Model Context Protocol (MCP; anthropicMCP2024) enables an LLM to interact with external systems and become more capable as an autonomous decision-making agent. CICERO (meta2022human) demonstrates strategic, human-level play with LLMs in the complex board game Diplomacy. Voyager (wang2023voyager) leverages Minecraft to illustrate the rising potential of LLMs as agents for open-ended exploration and skill acquisition. As LLM-agents become commonplace, new infrastructure is emerging to support agent-agent interaction, e.g. Google’s Agent2Agent protocol (A2A) enabling collaboration between LLM-based agents with varying capabilities, potentially from different organizations.
@@
Learning to resolve these scenarios is typically framed within multi-agent reinforcement learning (MARL). Social dilemmas are a specific subclass of MARL problems that are mixed-motive; neither fully cooperative nor fully competitive. Unlike single-agent RL, where an agent improves an objective in a static environment, in MARL each agent must adapt to the strategies of other agents, who can also be learning over time. This leads to non-stationarity, since the policy of each learning agent affects the collective outcome. Initial attempts using MARL to play social dilemmas were unsuccessful. Training agents based on small neural networks with naive MARL resulted in sub-optimal greedy strategies (sandholm1996multiagent). To address this, foerster2018learningopponentlearningawareness introduced Opponent Shaping (OS), an RL paradigm that explicitly considers agent interactions in hopes of steering their dynamics towards mutually beneficial outcomes. LOLA, the first OS algorithm, is capable of finding the pareto-optimal strategy of tit-for-tat in simple social dilemmas like the Iterated Prisoner’s Dilemma.
@@
To study this behavior, we introduce a novel testbed for social dilemmas in the LLM setting. The testbed includes small-scale social dilemma environments (duque2025advantagealignmentalgorithms) which we extend into the textual domain, as well as our new communication-based environment Trust and Split, designed to measure both cooperation and non-exploitability. Using this testbed, we conduct extensive experiments across a range of modern LLMs and find that naive MARL consistently produces greedy behavior across all environments. Probing further, we show that even state-of-the-art closed-source models are exploitable when facing agents trained with naive MARL. These results underscore that current LLMs are not yet prepared to robustly operate in real-world multi-agent settings and highlight a novel risk that arises from the consistent failure of naive MARL fine-tuning to produce cooperative and non-exploitable behavior in environments that include social dilemmas.
To overcome this issue, we adapt Advantage Alignment (duque_advantage_2025), a recent scalable opponent-learning awareness algorithm, to train LLM agents that cooperate reliably and resist exploitation in social dilemma environments. When trained with Advantage Alignment, we find that agents learn the non-exploitable and effective tit-for-tat strategy in the classic Iterated Prisoner’s Dilemma. In Trust and Split, Advantage Alignment agents learn to cooperate with cooperative players as well as themselves, while remaining robust against greedy players.
@@
  • Developing a social dilemma testbed tailored for LLMs, including standard environments and our novel Trust and Split environment, which requires communication to achieve high welfare.
  • Demonstrating that naive MARL leads to greedy, suboptimal agents across this testbed for a range of open-source LLMs, and showing that even state-of-the-art closed-source LLMs are vulnerable to exploitation by greedy RL-trained agents.
  • Adapting the Advantage Alignment algorithm (duque2025advantagealignmentalgorithms) to the LLM setting to train agents that reliably achieve cooperative, non-exploitable behavior across all environments in the testbed.
@@ 2.2 Multi-agent Reinforcement Learning
In a Markov Game, each agent $j$ attempts to maximize its objective. For each agent, the multi-agent state-value function is defined as $V^{j}(s):=\mathbb{E}_{\mathbf{a}\sim\Pi(s)}\left[r^{j}(s,\mathbf{a})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,\mathbf{a})}\left[V^{j}(s^{\prime})\right]\right]$, the action-value function as $Q^{j}(s,\mathbf{a}):=r^{j}(s,\mathbf{a})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,\mathbf{a})}\left[V^{j}(s^{\prime})\right]$, and the advantage function as $A^{j}(s,\mathbf{a}):=Q^{j}(s,\mathbf{a})-V^{j}(s)$. For any Markov Decision Process, the REINFORCE (williams_simple_1992) algorithm uses unbiased estimates of the gradient of the state-value function with respect to the parameters of $\pi$ in order to perform gradient ascent. RLOO (ahmadian2024basicsrevisitingreinforcestyle) reduces the variance of REINFORCE by introducing a simple baseline substraction. RLOO can easily be extended to the multi-agent case by independently updating each policy $j$ with $\nabla_{\theta^{j}}J^{j}(\Pi)=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\Pi}}\Bigl[\sum_{t=0}^{\infty}\gamma^{t}A^{j}(s_{t},\mathbf{a}_{t})\nabla_{\theta^{j}}\log\pi^{j}(\mathbf{a}_{t}^{j}|s_{t})\Bigr]$, the multi-agent advantage function being computed using an RLOO-style baseline (described in section 3). In the context of this paper, the naive MARL algorithm follows this formulation and is called multi-agent RLOO. We also consider the naive cooperative variant multi-agent RLOO with sum of rewards, which is algorithmically equivalent except for the fact that the reward functions of each agent are changed to $r(s,\mathbf{a}):=\sum_{j=1}^{n}r^{j}(s,\mathbf{a})$. That is, each agent optimizes for the sum of expected discounted returns across all agents. This formulation encourages agents to learn policies that maximize overall welfare rather than focusing on individual benefits.
@@ 2.3 Social Dilemmas
In a zero‑sum game, the agents’ payoffs always add up to zero; every gain for one side is matched by an equal loss for the other. Consequently, in a two-player zero‑sum setting, cooperation does not offer benefit. In this work, we focus on general‑sum games, where total payoffs are not fixed, and agents may improve their outcomes without necessarily diminishing those of others, thereby creating the possibility of mutually beneficial cooperation. More precisely, we focus on social dilemmas, general-sum games in which agents face a tension between their short-term individual benefit and long-term collective welfare. In these settings, each agent has a short-term incentive to act selfishly (i.e., not cooperate), but if all agents do so, the resulting outcome leads to reduced overall welfare,, i.e. a lower total sum of discounted returns for all agents. However, if an agent is unconditionally cooperative, other rational agents will exploit it and reduce its welfare to increase theirs. The focus of this paper is on a stronger alternative strategy, which incentivizes rational agents to behave in its best interest, achieving high collective welfare while avoiding exploitation.
@@ 2.4 Opponent Shaping
Prior work shows that small neural networks trained with naive MARL tend to converge to the Always Defect strategy in IPD (sandholm1996multiagent). More recently, foerster2018learningopponentlearningawareness demonstrated that this undesirable outcome also arises with policy gradient methods. These approaches assume that the environment is stationary, which is valid in single-agent setting, but not in multi-agent setting where other learning agents create non-stationarity. LOLA (foerster2018learningopponentlearningawareness) removed the assumption of a static environment in markov games and included a model of a learning agent in its update. By explicitly modeling how opponent learning is affected by an agent’s action, LOLA was able to learn the tit-for-tat strategy in IPD. Unfortunately, LOLA’s computational complexity is quadratic in the number of parameters of the agent, making it impractical for LLMs.
Advantage Alignment (duque2025advantagealignmentalgorithms) is an opponent-shaping algorithm that instead focuses on the Q-values of both the agent and its opponent. Assuming that agents act proportionally to the exponent of their Q-value, Advantage Alignment aims to align an opponent’s Q-value with its own. This leads to a simple modification to the advantages used in the policy gradient term of a REINFORCE estimator. Advantage Aligment has been shown to solve social dilemmas in scenarios with high dimensional state representations (e.g. pixel spaces), partial observability, and continuous action spaces. Given its performance in complex scenarios, we chose Advantage Alignment as a prime candidate to train LLMs to find cooperative and non-exploitable strategies.
3 Method
@@
Estimating advantages with value networks has proven challenging in the context of LLM training, often leading to unstable or ineffective results (kazemnejad2024vineppo). Recent work such as RLOO (ahmadian2024basicsrevisitingreinforcestyle) and GRPO (shao2024deepseekmath) has shown that baseline-based approaches provide more stable and efficient advantage estimates. These approaches sample multiple trajectories for a given prefix, and compute the advantage for each trajectory as the difference between its discounted return and the mean discounted return of the remaining trajectories. However, scaling this approach to multi-round, multi-agent settings is infeasible because the number of trajectories needed grows exponentially. In our experiments, we build on these ideas and extend them to multi-agent LLM training. We divide each batch of rollouts into $k$ common random number (CRN) groups, each of which uses a fixed random seed to generate the environment stochasticity. This ensures that, within a CRN group, the variance in discounted returns comes only from the agent’s actions and not from the environment. This is similar in spirit to GRPO and RLOO, except trajectories share a fixed environment context rather than a shared prefix. Particularly, let $A^{i}(s_{t},a_{t})$ denote the advantage for agent $i$. We estimate it using a leave-one-out group baseline computed over the $k$ games of its CRN group at each time step $t$: $G(a_{t}^{(i)},s_{t})-\frac{1}{k-1}\sum_{j\neq i}G(a_{t}^{(j)},s_{t})$ where $G(a_{t}^{(i)},s_{t})$ is the discounted return for action $a_{t}^{(i)}$ taken in state $s_{t}$. This group-relative baseline avoids the need for a learned value function, simplifies advantage computation, and enables multi-turn RL with LLMs in our multi-agent settings. We refer to this algorithm as multi-agent RLOO in the rest of the paper.
@@
Each agent’s policy $\pi_{i}$ is parameterized by $\theta_{i}$ and implemented via LoRA finetuning (hu2022lora). Throughout our experiments, we refer to the first player as Alice and the second as Bob. We use self-play, i.e, the same set of parameters for both agents, conditioned on different game contexts based on their roles. This ensures that memory usage doesn’t scale with the number of agents and the model size we used is sufficient enough to handle the complexity of the different roles. Maintaining opponent diversity is essential for self-play, and it is particularly important in social dilemmas, where defection equilibria can trap learning. Without diversity, exploration suffers and agents may remain stuck in defecting strategies. Following duque2025advantagealignmentalgorithms, we preserve opponent diversity through a agent buffer that stores earlier versions of the self-play agent. This is straightforward to implement because each agent is represented by a LoRA checkpoint, roughly $0.1\%$ of the model parameters, which can be saved and reloaded with minimal overhead. For each game, with probability $\rho$, the opponent is sampled from the agent buffer. With probability $1-\rho$, the opponent is simply the current version of the agent using the latest LoRA parameters. We use $\rho=1/2$ as the default setting, and it works well in our experiments.
@@ 4 Social Dilemma Testbed
Figure 1: One round of Trust and Split. Each player receives a private rock-paper-scissors hand that determines how much they value the coins, sends one message in turn, and then submits a proposal. Payoffs follow the split rule. Both hands and proposals are revealed before the next round starts.
In this section, we study the behavior of LLM agents trained with naive MARL in social dilemma environments. To support this, we develop a novel testbed tailored for LLMs to evaluate the effects of MARL training on cooperation and resistance to exploitation. An exact description of the game prompt in all the environments is provided in the Appendix 12.
@@
IPD is a two-player game where agents repeatedly and simultaneously choose to either Cooperate (C) or Defect (D) in each round. The per‑round pay‑off matrix used in our experiments is provided in Table 11 in the Appendix . We include IPD in our testbed because it is one of the most widely studied social dilemmas. However, since it is also likely presented in the training data of LLMs, we obfuscate the nature of the game by removing any mention of “Prisoner’s Dilemma” and replace the action labels from Cooperate and Defect to A and B, respectively. This allows us to test how well LLMs generalize beyond memorization and to examine how RL interacts with any prior knowledge the model may have about this social dilemma.
@@
This environment is textual version of the negotiation game used in duque_advantage_2025. In this game, there are three item categories (hats, books and balls) to split at each round. The values of each item are public for both the agents. At each round, item values are sampled as follows: (1) each item category is assigned a value of either $1$ or $10$ at random, (2) at least one item category must have different values for the two agents, creating a conflict and a social dilemma, and (3) the total value across all items is the same for both the agents in that round. Proposals and payoffs are revealed after the end of each round. This variant supports reciprocity without the need for communication. The split rule (proposal mechanism) from the Negotiation Game (Cao2018EmergentCT; duque2025advantagealignmentalgorithms), providing a better learning signal for training agents in this dilemma. More precisely, let $p_{k,a}$ be the proposal for the $k$’th item category from agent $a$ and $q_{k}$ be the quantity available. The allocation received by agent $a$ is $q_{k,a}=q_{k}p_{k,a}/\max\bigl(q_{k},p_{k,a}+p_{k,b}\bigr)$ and similarly for agent $b$. The resulting payoffs are $v_{a}\times q_{k,a}$ and $v_{b}\times q_{k,b}$, respectively. This particular mechanism‑design choice removes the need for explicit agreement and provides a training signal in each round.
Trust and Split
@@
While IPD and Split No-Comm captures the fundamental dilemma, it lacks the richness of real-world strategic interactions. To address this, we propose Trust and Split, a novel environment that builds on Split No-Comm by adding communication. However, communication for negotiating multiple items leads to long contexts that can be challenging for opensource LLMs. liao_efficacy_2024 find that LLMs upto the scale of 7070B struggle to follow instruction in multiple item setting across multiple rounds. Trust and Split handles this limitation using a single item: coins. It is designed so that effective performance still requires communication. The setup allows for variety of behaviors, including bluffing, exaggeration, and cooperative negotiation. A visualization of a round in this environment is detailed in figure 1. At the beginning of each round, each player is assigned an exclusive private hand among {rock, paper, scissors}. The agent with the lower hand values each coin at $1$, while the agent with upper hand values each coin at $10$. Since neither player knows the other’s hand, they are incentivized to communicate to infer values and play effectively. Each agent can then negotiate with the other agent, one message at a time. We currently limit the number of messages to one per agent to ensure that we can train these agents across multiple rounds. After the messaging phase, both agents submit their proposals simultaneously. They then receive their payoffs based on their coin values and the quantities allocated by the split rule. Before continuing to the next round, the hands and proposals are revealed to both players, allowing reciprocity. In this environment, the starting agent alternates every round, and hands are assigned so that, in expectation, both agents receive an equal number of upper hands. The strategy that maximizes payoffs for both the agents is to truthfully communicate hands and allocate all items to the agent who values them more in each round, while remaining non-exploitable.

### v2

@@ Abstract
As agentic AI becomes more widespread, agents with distinct and possibly conflicting goals will interact in complex ways. These multi-agent interactions pose a fundamental challenge, particularly in social dilemmas, where agents’ individual incentives can undermine collective welfare. While reinforcement learning (RL) has been effective for aligning large language models (LLMs) in the single-agent regime, prior small-network results suggest that standard RL in multi-agent settings often converges to defecting, self-interested policies. We show the same effect in LLMs: despite cooperative priors, RL-trained LLM agents develop opportunistic behavior that can exploit even advanced closed-source models. To address this tendency of RL to converge to poor equilibria, we adapt a recent opponent-learning awareness algorithm, Advantage Alignment, to fine-tune LLMs toward multi-agent cooperation and non-exploitability. We then introduce a group-relative baseline that simplifies advantage computation in iterated games, enabling multi-agent training at LLM scale. We also contribute a novel social dilemma environment, Trust-and-Split, which requires natural language communication to achieve high collective welfare. Across a wide range of social dilemmas, policies learned with Advantage Alignment achieve higher collective payoffs while remaining robust against exploitation by greedy agents. We release all of our code to support future work on multi-agent RL training for LLMs.  [^1] https://github.com/dereckpiche/AdAlignLLM
@@ 1 Introduction
LLMs undergo large-scale pretraining, instruction tuning, and reinforcement learning, and continue to exhibit increasingly advanced capabilities (deepseekr1). Coupled with decreasing deployment costs and improved adaptability to downstream tasks, these trends enhance the commercial and practical viability of LLM agents across a wide range of applications. Recent efforts are already translating this potential into concrete systems. Anthropic’s Model Context Protocol (MCP; anthropicMCP2024) enables an LLM to interact with external systems and become more capable as an autonomous decision-making agent. At the same time, LLM agents are now being deployed in real applications, from code generation and software development assistance (chen2021evaluating) to e-commerce transactions and personalized information curation (openai2024agents; openai2024pulse). New infrastructure is also emerging to support agent-agent interaction, such as Google’s Agent2Agent protocol (A2A), which enables collaboration among LLM-based agents with varying capabilities, potentially across different organizations.
@@
Learning to resolve these scenarios is typically framed within multi-agent reinforcement learning (MARL). Social dilemmas are a specific subclass of MARL problems that are mixed-motive; neither fully cooperative nor fully competitive. Unlike single-agent RL, where an agent improves an objective in a static environment, in MARL each agent must adapt to the strategies of other agents, who can also be learning over time. This leads to non-stationarity, since the policy of each learning agent affects the collective outcome. Initial attempts using MARL to play social dilemmas were unsuccessful. Training agents based on small neural networks with naive MARL resulted in sub-optimal greedy strategies (sandholm1996multiagent). To address this, foerster2018learningopponentlearningawareness introduced Opponent Shaping (OS), an RL paradigm that explicitly considers agent interactions in hopes of steering their dynamics toward mutually beneficial outcomes. LOLA, the first OS algorithm, is capable of finding the pareto-optimal strategy of tit-for-tat in simple social dilemmas like the Iterated Prisoner’s Dilemma.
@@
To study this behavior, we introduce a novel testbed for social dilemmas in the LLM setting. The testbed includes small-scale social dilemma environments (duque2025advantagealignmentalgorithms) which we extend to the textual domain, as well as our new communication-based Trust-and-Split environment, designed to measure both cooperation and non-exploitability. Using this testbed, we conduct extensive experiments across a range of modern LLMs and find that naive MARL consistently produces greedy behavior across all environments. Probing further, we show that even state-of-the-art closed-source models are exploitable when facing agents trained with naive MARL. These results underscore that current LLMs are not yet prepared to operate robustly in real-world multi-agent settings. Together, they provide a novel insight: failure in multi-agent settings can arise simply from naive MARL fine-tuning in a social dilemma.
To overcome this issue, we adapt Advantage Alignment (duque_advantage_2025), a recent opponent shaping algorithm, to train LLM agents that cooperate reliably and resist exploitation in social dilemma environments. Specifically, we introduce a group-relative baseline to compute advantages in multi-round settings and implement an agent buffer with LoRA (hu2022lora) to maintain diversity during training. When trained with Advantage Alignment using these design choices, we find that agents learn the non-exploitable and effective tit-for-tat strategy in the classic Iterated Prisoner’s Dilemma. In complex environments like Split No-Comm and Trust-and-Split, Advantage Alignment agents learn to cooperate with cooperative players as well as themselves, while remaining robust against greedy players.
@@
  • Developing a social dilemma testbed for LLMs, including standard environments and our novel Trust-and-Split environment, where achieving high welfare requires communication.
  • Demonstrating that naive MARL leads to greedy, suboptimal agents across this testbed for a range of open-source LLMs, and that even state-of-the-art closed-source LLMs are vulnerable to exploitation by greedy RL-trained agents.
  • Adapting the Advantage Alignment algorithm (duque2025advantagealignmentalgorithms) to the LLM setting to train agents that achieve cooperative, non-exploitable behavior across this testbed.
@@ 2.2 Multi-agent Reinforcement Learning
In a Markov Game, each agent $j$ attempts to maximize its objective. For each agent, the multi-agent state-value function is defined as $V^{j}(s):=\mathbb{E}_{\mathbf{a}\sim\Pi(s)}\left[r^{j}(s,\mathbf{a})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,\mathbf{a})}\left[V^{j}(s^{\prime})\right]\right]$, the action-value function as $Q^{j}(s,\mathbf{a}):=r^{j}(s,\mathbf{a})+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,\mathbf{a})}\left[V^{j}(s^{\prime})\right]$, and the advantage function as $A^{j}(s,\mathbf{a}):=Q^{j}(s,\mathbf{a})-V^{j}(s)$. For any Markov Decision Process, the REINFORCE (williams_simple_1992) algorithm uses unbiased estimates of the gradient of the state-value function with respect to the parameters of $\pi$ in order to perform gradient ascent. GRPO (shao2024deepseekmath) reduces the variance of REINFORCE by introducing a simple baseline subtraction. GRPO can easily be extended to the multi-agent case by independently updating each policy $j$ with $\nabla_{\theta^{j}}J^{j}(\Pi)=\mathbb{E}_{\tau\sim\text{Pr}_{\mu}^{\Pi}}\Bigl[\sum_{t=0}^{\infty}\gamma^{t}A^{j}(s_{t},\mathbf{a}_{t})\nabla_{\theta^{j}}\log\pi^{j}(\mathbf{a}_{t}^{j}|s_{t})\Bigr]$, the multi-agent advantage function being computed using a GRPO-style baseline (described in section 3). In the context of this paper, the naive MARL algorithm follows this formulation and is called multi-agent GRPO. We also consider the naive cooperative variant multi-agent GRPO with sum of rewards, which is algorithmically equivalent except for the fact that the reward functions of each agent are changed to $r(s,\mathbf{a}):=\sum_{j=1}^{n}r^{j}(s,\mathbf{a})$. That is, each agent optimizes for the sum of expected discounted returns across all agents. This formulation encourages agents to learn policies that maximize overall welfare rather than focusing on individual benefits.
@@ 2.3 Social Dilemmas
In a zero‑sum game, the agents’ payoffs always add up to zero; every gain for one side is matched by an equal loss for the other. Consequently, in a two-player zero‑sum setting, cooperation does not offer benefit. In this work, we focus on general‑sum games, where total payoffs are not fixed, and agents may improve their outcomes without necessarily diminishing those of others, thereby creating the possibility of mutually beneficial cooperation. More precisely, we focus on social dilemmas, general-sum games in which agents face a tension between their short-term individual benefit and long-term collective welfare. In these settings, each agent has a short-term incentive to act selfishly (i.e., not cooperate), but if all agents do so, the resulting outcome leads to reduced overall welfare, i.e. a lower total sum of discounted returns for all agents. However, if an agent is unconditionally cooperative, other rational agents will exploit it and reduce its welfare to increase theirs. The focus of this paper is on a stronger alternative strategy, which incentivizes rational agents to behave in its best interest, achieving high collective welfare while avoiding exploitation.
@@ 2.4 Opponent Shaping
Prior work shows that small neural networks trained with naive MARL tend to converge to the Always Defect strategy in IPD (sandholm1996multiagent). More recently, foerster2018learningopponentlearningawareness demonstrated that this undesirable outcome also arises with policy gradient methods. These approaches assume that the environment is stationary, which is valid in a single-agent setting, but not in a multi-agent setting where other learning agents create non-stationarity. LOLA (foerster2018learningopponentlearningawareness) removed the assumption of a static environment in markov games and included a model of a learning agent in its update. By explicitly modeling how opponent learning is affected by an agent’s action, LOLA was able to learn the tit-for-tat strategy in IPD. Unfortunately, LOLA’s computational complexity is quadratic in the number of parameters of the agent, making it impractical for LLMs.
Advantage Alignment (duque2025advantagealignmentalgorithms) is an opponent-shaping algorithm that instead focuses on the Q-values of both the agent and its opponent. Assuming that agents act proportionally to the exponent of their Q-value, Advantage Alignment aims to align an opponent’s Q-value with its own. This leads to a simple modification to the advantages used in the policy gradient term of a REINFORCE estimator. Advantage Alignment has been shown to solve social dilemmas in scenarios with high dimensional state representations (e.g. pixel spaces), partial observability, and continuous action spaces. Given its performance in complex scenarios, we chose Advantage Alignment as a prime candidate to train LLMs to find cooperative and non-exploitable strategies.
3 Advantage Alignment for LLMs
@@
Estimating advantages with value networks has proven challenging in the context of LLM training, often leading to unstable or ineffective results (kazemnejad2024vineppo). Recent work such as RLOO (ahmadian2024basicsrevisitingreinforcestyle) and GRPO (shao2024deepseekmath) has shown that baseline-based approaches provide more stable and efficient advantage estimates. These approaches sample multiple trajectories for a given prefix, and compute the advantage for each trajectory as the difference between its discounted return and the mean discounted return of the remaining trajectories. However, scaling this approach to multi-round, multi-agent settings is infeasible because the number of trajectories needed grows exponentially. In our experiments, we build on these ideas and extend them to multi-agent LLM training. We divide each batch of rollouts into $k$ common random number (CRN) groups, each of which uses a fixed random seed to generate the environment stochasticity. This ensures that, within a CRN group, the variance in discounted returns comes only from the agent’s actions and not from the environment. This is similar in spirit to GRPO and RLOO, except trajectories share a fixed environment context rather than a shared prefix. In particular, let $A^{i}(s_{t},a_{t})$ denote the advantage for agent $i$. We estimate it using a leave-one-out group baseline computed over the $k$ games of its CRN group at each time step $t$: $G(a_{t}^{(i)},s_{t})-\frac{1}{k-1}\sum_{j\neq i}G(a_{t}^{(j)},s_{t})$ where $G(a_{t}^{(i)},s_{t})$ is the discounted return for action $a_{t}^{(i)}$ taken in state $s_{t}$. This group-relative baseline avoids the need for a learned value function, simplifies advantage computation, and enables multi-turn RL with LLMs in our multi-agent settings. We refer to this algorithm as multi-agent GRPO in the rest of the paper.
@@
Each agent’s policy $\pi_{i}$ is parameterized by $\theta_{i}$ and implemented via LoRA finetuning (hu2022lora). Throughout our experiments, we refer to the first player as Alice and the second as Bob. We use self-play, i.e., the same set of parameters for both agents, conditioned on different game contexts based on their roles. This ensures that memory usage doesn’t scale with the number of agents and the model size we used is sufficient to handle the complexity of the different roles. Maintaining opponent diversity is essential for self-play, and it is particularly important in social dilemmas, where defection equilibria can trap learning. Without diversity, exploration suffers and agents may remain stuck in defecting strategies. Following duque2025advantagealignmentalgorithms, we preserve opponent diversity through an agent buffer that stores earlier versions of the self-play agent. This is straightforward to implement because each agent is represented by a LoRA checkpoint, roughly $0.1\%$ of the model parameters, which can be saved and reloaded with minimal overhead. For each game, with probability $\rho$, the opponent is sampled from the agent buffer. With probability $1-\rho$, the opponent is simply the current version of the agent using the latest LoRA parameters. We use $\rho=1/2$ as the default setting, and it works well in our experiments. For both multi-agent GRPO and its sum-of-rewards variant, the agent buffer made no noticeable difference, and for computational reasons we did not apply it to these methods.
@@ 4 Social Dilemma Testbed
Figure 1: One round of Trust-and-Split. Each player receives a private rock-paper-scissors hand that determines how much they value the coins, sends one message in turn, and then submits a proposal. Payoffs follow the split rule. Both hands and proposals are revealed before the next round starts.
In this section, we study the behavior of LLM agents trained with naive MARL in social dilemma environments. To support this, we develop a novel testbed tailored for LLMs to evaluate the effects of MARL training on cooperation and resistance to exploitation. An exact description of the game prompt in all the environments is provided in Appendix 13.
@@
IPD is a two-player game where agents repeatedly and simultaneously choose to either Cooperate (C) or Defect (D) in each round. The per‑round pay‑off matrix used in our experiments is provided in Table 12 in the appendix. We include IPD in our testbed because it is one of the most widely studied social dilemmas. However, since it is also likely presented in the training data of LLMs, we obfuscate the nature of the game by removing any mention of “Prisoner’s Dilemma” and replace the action labels Cooperate and Defect with A and B, respectively. This allows us to test how well LLMs generalize beyond memorization and to examine how RL interacts with any prior knowledge the model may have about this social dilemma.
@@
This environment is a textual version of the negotiation game used in duque_advantage_2025. In this game, there are three item categories (hats, books and balls) to split at each round. The values of each item are public for both agents. At each round, item values are sampled as follows: (1) each item category is assigned a value of either $1$ or $10$ at random, (2) at least one item category must have different values for the two agents, creating a conflict and a social dilemma, and (3) the total value across all items is the same for both agents in that round. Proposals and payoffs are revealed after the end of each round. This variant supports reciprocity without the need for communication. The split rule (proposal mechanism) from the Negotiation Game (Cao2018EmergentCT; duque2025advantagealignmentalgorithms), provides a better learning signal for training agents in this dilemma. More precisely, let $p_{k,a}$ be the proposal for the $k$’th item category from agent $a$ and $q_{k}$ be the quantity available. The allocation received by agent $a$ is $q_{k,a}=q_{k}p_{k,a}/\max\bigl(q_{k},p_{k,a}+p_{k,b}\bigr)$ and similarly for agent $b$. The resulting payoffs are $v_{a}\times q_{k,a}$ and $v_{b}\times q_{k,b}$ respectively. This particular choice removes the need for explicit agreement and ensures that both agents receive a learning signal every round.
Trust-and-Split
@@
While IPD and Split No-Comm capture the fundamental dilemma, they lack the richness of real-world strategic interactions. Existing negotiation environments involve longer interactions (davidson2024evaluating; lewis_deal_2017), which make them less feasible to train and more difficult for characterizing robust strategies that maximize collective payoff. Moreover, liao_efficacy_2024 find that LLMs up to the scale of 7070B struggle to follow instructions in multi-item settings across multiple rounds. To address these limitations, we propose Trust-and-Split, a novel environment that builds on Split No-Comm by adding communication. Trust-and-Split uses a single item, coins, which avoids the complexity of multi-item negotiation while still requiring communication for effective performance. A visualization of a round in this environment is detailed in Figure 1. At the beginning of each round, each player is assigned an exclusive private hand among {rock, paper, scissors}. The agent with the lower hand values each coin at $1$, while the agent with the upper hand values each coin at $10$. Since neither player knows the other’s hand, they are incentivized to communicate to infer values and play effectively. Each agent can then negotiate with the other agent, one message at a time. We currently limit the number of messages to one per agent to ensure that we can train these agents across multiple rounds. The setup also allows for a variety of behaviors, including bluffing, exaggeration, and cooperative negotiation. After the messaging phase, both agents submit their proposals simultaneously. They then receive their payoffs based on their coin values and the quantities allocated by the split rule. Before continuing to the next round, the hands and proposals are revealed to both players, allowing reciprocity. In this environment, the starting agent alternates every round, and hands are assigned so that in expectation, both agents receive an equal number of upper hands. The strategy that maximizes payoffs for both agents is to truthfully communicate hands and allocate all items to the agent who values them more in each round, while remaining non-exploitable.